OpenStudy (anonymous):

what is the slant asymptote of f(x)=(4x^3+3x^2-1)/(2x^2-x)?

sam (.sam.):

\[\frac{4 x^3+3 x^2-1}{2 x^2-x}\to \pm \infty \text{ as }x\to 0\] \[\frac{4 x^3+3 x^2-1}{2 x^2-x}\to \pm \infty \text{ as }x\to \frac{1}{2}\]

OpenStudy (anonymous):

to find the slant asymptote, divide

OpenStudy (anonymous):

same bike

OpenStudy (anonymous):

i am a trend setter

OpenStudy (anonymous):

get \[2x+\frac{5}{2}+\text{remainder}\] ignore the remainder